OpenStudy (anonymous):

Use the quadratic formula to find any x-intercepts on the graph of the equation. y=4x^2-2x-5

OpenStudy (anonymous):

what are you having trouble with on this?

OpenStudy (anonymous):

@HawkCrimson

OpenStudy (anonymous):

\[(-b \sqrt[\pm]{b ^{2}-4ac} )/2a\]

OpenStudy (anonymous):

where ax^2 +bx + c

OpenStudy (anonymous):

okay your first equation was right!

OpenStudy (anonymous):

just plug in the values and work it out :)

OpenStudy (anonymous):

Trying to do that now =l

OpenStudy (anonymous):

so I got, \[2\pm \sqrt{-76} / -4 \] doesnt seem right?

OpenStudy (anonymous):

u made a mistake in the square root and the -4

OpenStudy (anonymous):

ill show u know

OpenStudy (anonymous):

I got it further down to \[-2\pm 2\sqrt{38}/ -4 \]

OpenStudy (anonymous):

\[b ^{2}-4ac\] (-2)^2 - 4(4)(-5) 4 -4(-20) 4+80 = 84

OpenStudy (anonymous):

2a 2(4) 8 not -4

OpenStudy (anonymous):

\[2\sqrt[\pm]{84}/8\]

OpenStudy (anonymous):

understand?

OpenStudy (anonymous):

yeah that makes sense, now it's factoring it down.

OpenStudy (anonymous):

no, next u square root the 84

OpenStudy (anonymous):

so u'll get 9.17

OpenStudy (anonymous):

\(84=4\times 21\) so \(\sqrt{84}=\sqrt{4}\sqrt{21}=2\sqrt{21}\)

OpenStudy (anonymous):

you get \[\frac{2\pm2\sqrt{21}}{8}=\frac{2(1\pm\sqrt{21})}{8}=\frac{1\pm\sqrt{21}}{4}\]

OpenStudy (anonymous):

umm @satellite73 that is all fun and gud but this question isnt so complicated..

OpenStudy (anonymous):

it is if you want to write your answer in simplest radical form if you want a decimal, then you use a calculator

OpenStudy (anonymous):

well thanks i didnt know that :) though we dont need to do that ill keep that in mind thx :D

OpenStudy (anonymous):

He did have the correct answer.

OpenStudy (anonymous):

Use the quadratic formula to find any x-intercepts on the graph of the equation. y=4x^2-2x-5 ^ That didnt ask for it in the racinal form but ok xD

OpenStudy (anonymous):

radical*

OpenStudy (anonymous):

that is true, but that's how the instructor had the answer =) no biggie. it wasn't specific enough.

OpenStudy (anonymous):

lol ok but on the record, i did most of the work XP